# Real-Time Dynamics in Two Dimensions with Tensor Network States via Time-Dependent Variational Monte Carlo

### v1

@@ Abstract
Reliably simulating two-dimensional many-body quantum dynamics with projected entangled pair states (PEPS) has long been a difficult challenge. In this work, we overcome this barrier for low-energy quantum dynamics by developing a stable and efficient time-dependent variational Monte Carlo (tVMC) framework for PEPS. By analytically removing all gauge redundancies of the PEPS manifold and exploiting tensor locality, we obtain a numerically well-conditioned stochastic reconfiguration (SR) equation amenable to robust solution using the efficient Cholesky decomposition, enabling long-time evolution in previously inaccessible regimes. We demonstrate the power and generality of the method through four representative real-time problems in two dimensions: (I) chiral edge propagation in a free-fermion Chern insulator; (II) fractionalized charge transport in a fractional Chern insulator; (III) vison confinement dynamics in the Higgs phase of a $\mathbb{Z}_{2}$ lattice gauge theory; and (IV) superfluidity and critical velocity in interacting bosons. All simulations are performed on $12\times 12$ or $13\times 13$ lattices with evolution times $T=10$ to $12$ using modest computational resources — 1 to 5 days on a single GPU card. Where exact benchmarks exist (case I), PEPS–tVMC matches free-fermion dynamics with high accuracy up to $T=12$. These results establish PEPS–tVMC as a practical and versatile tool for real-time quantum dynamics in two dimensions. The method significantly extends the reach of classical tensor-network simulations for studying elementary excitations in quantum many-body systems and provides a valuable computational counterpart to emerging quantum simulators.
@@ I Introduction
Projected entangled pair states (PEPS) [19] provide a complementary variational family whose locality and multi-linearity make them particularly appealing for tVMC. Within VMC, only a single layer of PEPS tensors needs to be contracted, making the method substantially more efficient than conventional PEPS update schemes [20, 21, 22]. Furthermore, VMC allows parallel accumulation of samples, making the approach well suited for modern GPU architectures.
@@
While dynamical simulations with tensor network states are enormously successful in 1D [27, 28, 29, 30], the situation in 2D is much more difficult. Traditional approaches to dynamical PEPS have relied on full-update methods [20, 21, 22], which face several challenges: (a) high cost of double layer contraction, (b) difficulty in fixing the PEPS gauge freedom, and (c) loss of positivity in the norm matrix approximately obtained from the contraction of the double-layer environment. We show that all three issues are naturally resolved within PEPS–tVMC at the expense of manageable statistical noise.
@@ III Stochastic reconfiguration with PEPS
In this section we derive the SR equations for PEPS, analyze the structure of their null space, and describe how gauge redundancy can be eliminated efficiently in closed form.
Figure 1: PEPS structure and gauge freedom. (a) A $3\times 3$ PEPS network with local tensors AA. (b) A single tensor AA with physical index $p$ and virtual indices $(l,r,d,u)$. (c) Local gauge transformation on a virtual bond: inserting an invertible matrix $X$ and its inverse $X^{-1}$ between two adjacent tensors leaves the global PEPS wavefunction invariant.
@@
Removal of gauge redundancy.
We now outline the procedure for eliminating null vectors prior to solving the SR equation. First, we construct a projection matrix $P$ onto the subspace orthogonal to the constraint vectors $\bm{\mathbf{w}}$, thereby reducing the number of independent gauge vectors to
@@
III.1 Gauge-fixing with minSR
@@
A key observation is that the PEPS gauge redundancy does not appear at the minSR level, because the parameter index is contracted away, thus gauge-fixing is automatically achieved. Assuming decorrelation between samples, the matrix $\mathcal{O}^{\dagger}$ has a single null vector $1,1,\dots,1]^{T}$, which we remove with the same QR projection described above. More generally, the null vector of $\mathcal{O}^{\dagger}$ is the weight vector $W$ used in performing the statistical average: $\braket{A}\approx\sum_{s=1}^{N_{s}}W_{s}A_{s}/N_{s}$, which in our case is simply the identity weight. A non-trivial weight vector would arise if reweighing is used in the Monte Carlo estimate.
III.2 Small-$o$ trick in minSR
@@
from which the full minSR matrix $\mathcal{O}\mathcal{O}^{\dagger}$ can be reconstructed with the memory cost of $o$ instead of $O$ (see the Supplementary Material [35] for the explicit construction).
@@
III.3 Comparison with double-layer methods
We briefly discuss the two difficulties of double-layer methods mentioned in the introduction: gauge fixing and positivity of the norm matrix.
@@ III.3 Comparison with double-layer methods
Unlike matrix product states in 1D [1], PEPS does not possess a canonical form that fixes the gauge freedoms. Ref. [22] introduces local gauge fixing for full-update gate applications, but no globally consistent gauge-fixing procedure is known. In fact, as an alternative, a variational subfamily — isometric PEPS [36, 37, 38] — is introduced to explicitly preserve the isometric gauge at the expense of representative expressiveness. However, the necessary shifting of isometric centers introduces significant errors and leads to suboptimal results. In contrast, PEPS–tVMC achieves global gauge fixing efficiently through a single QR projection or simply using the minSR formulation, without any loss of representational power.
@@
Because PEPS $\psi$ is an unnormalized ansatz, to perform a variational update, one needs to also form the norm matrix $N$ accounting for the norm $\braket{\psi|\psi}$. In practice, approximate contraction leads to violations of positivity of $N$, which destabilizes the generalized eigenvalue problems used in variational updates, especially in time evolution. tVMC avoids this issue entirely: normalization is enforced simply by subtracting $O_{\rm avg}$ and $h_{\rm avg}$ in Eq. 5, requiring no explicit norm matrix.
@@ IV Results
We now apply PEPS–tVMC to study 4 distinct two-dimensional models. These include both bosonic and fermionic systems, free and interacting dynamics, and examples with gauge symmetry or nontrivial topology. We use PEPS sequencial sampling [39, 40] to obtain the samples. The simulation details and hyperparameters of the tVMC are presented in the Supplementary Material.
@@ IV.1 Free Fermion Chern insulator
As shown in Fig. 2, the excitation propagates unidirectionally along the edge. PEPS–tVMC faithfully reproduces the edge motion, its speed, and its spatial profile, in excellent agreement with the exact free-fermion numerics. We emphasize that despite the model being noninteracting, its entanglement is significant and the dynamics is highly non-trivial, making this a stringent benchmark for PEPS, especially at this lattice size and time scale.
@@ IV.3 $\mathbb{Z}_{2}$ lattice gauge theory with Higgs field
Next, using the gauge-invariant PEPS-VMC framework recently introduced by us [31], we examine the confinement dynamics in the $\mathbb{Z}_{2}$ gauge theory coupled to a Higgs field,
@@
subject to the Gauss-law constraint $\sigma_{i}^{z}\prod_{l\in i}Z_{l}=1$. $X_{l},Z_{l}$ are the Pauli $x,z$ matrices on the link $l$, representing the gauge field. $\sigma^{x}_{i},\sigma^{z}_{i}$ are the Pauli matrices on the site $i$, representing the Higgs field. $l_{\pm}$ are the sites at the ends of link $l$. Here $B_{p}=\prod_{l\in p}X_{l}$ measures the magnetic flux through plaquette $p$, and a violation of $B_{p}=1$ defines a vison excitation. The system has three ground state phases: confined (small $J$ and large $g$), Higgs (large $J$ and small $g$), and deconfined (small $J$ and small $g$) phases [52]. In the Higgs phase, the visons are expected to have a linearly attractive interaction, i.e. vison confinement. In the deconfined phase, the visons are expected to propagate freely, i.e. vison deconfinement.
@@
Recently, considerable progress has been made in implementing real-time gauge-theory dynamics on quantum devices [53, 54, 55]. Our results highlight that classical tensor-network approaches can serve as a highly effective and complementary platform for these studies. It would not be impossible for PEPS–tVMC to probe the intricate dynamics of glueballs in the (2+1)D Yang-Mills theory, after the gauge-invariant PEPS-VMC [31] is extended to the non-Abelian case.
@@ IV.4 Superfluid bosons
Figure 5: Real-time density response of a 2D bosonic superfluid to a moving obstacle. Snapshots show the deviation of the local density from 0.5 on a $13\times 13$ lattice for different times $T$ as the obstacle (red dot) is dragged along the $x$ direction. (a) $|\bm{\mathbf{v}}|=0.5$. (b) $|\bm{\mathbf{v}}|=5.0$.
@@ IV.5 Error in solving the SR equation
An important factor underlying the accuracy and efficiency of PEPS–tVMC is the numerical stability with which the SR equation—or its reduced minSR variant—can be solved. All simulations in this work solve the SR equation using a Cholesky decomposition of the SR matrix with a diagonal regulator of magnitude $10^{-8}$. Compared to the rank-revealing singular value decomposition (SVD) solver, the Cholesky approach is substantially faster and more parallelizable on GPUs, but it lacks the robustness of SVD when applied to ill-conditioned matrices. Recent NQS–tVMC studies often resort to the SVD solver for this reason [13, 16, 18], trading computational efficiency for numerical robustness. The fact that Cholesky remains stable in all simulations indicates that once gauge redundancies are analytically removed, the PEPS variational manifold becomes exceptionally stable for real-time local quench dynamics.
@@ V Discussion and Outlook
The examples explored in Sec. IV demonstrate the breadth of physical phenomena accessible with this approach. These results establish PEPS–tVMC as a general-purpose tool for simulating elementary excitations, transport phenomena, and real-time response in a wide range of quantum systems with realistic computational resources. The improvements developed here apply equally to ground-state PEPS-VMC calculations [59, 60, 61, 62]. Extensions to excited states, finite temperature [63], and open quantum systems are also promising avenues. In the case of finite temperature, if purification-based ansatz is used, the small-$o$ trick can save the memory cost by a factor of 16 for spinful fermions!
@@ Acknowledgements.
Part of the algorithm is tested using the TenPy code base [64, 65]. The code for production is implemented with JAX [66]. Y.W. is supported by a startup grant from the IOP-CAS. I thank my colleagues Yu-Xiang Zhang, Shi-Xin Zhang, and Tao Xiang for sharing computational resources with me. I benefited greatly from discussion and collaboration on various related projects with Bartholomew Andrews, Zhehao Dai, Jannes Nys, Sheng-Hsuan Lin, Sajant Anand, Michael Zaletel, Wen-Yuan Liu, Jiahang Hu, Jiajun Yu, and Yuntian Gu.
@@
    2def get_OOdag(jnp.ndarray: o, jnp.ndarray: p, int: K):
    3 r’’

### v2

@@ Abstract
Reliably simulating two-dimensional many-body quantum dynamics with projected entangled pair states (PEPS) has long been a difficult challenge. In this work, we overcome this barrier for low-energy quantum dynamics by developing a stable and efficient time-dependent variational Monte Carlo (tVMC) framework for PEPS. By analytically removing all gauge redundancies of the PEPS manifold and exploiting tensor locality, we obtain a numerically well-conditioned stochastic reconfiguration equation amenable to robust solution using the efficient Cholesky decomposition, enabling long-time evolution in previously inaccessible regimes. We explain how the difficulties associated with the traditional PEPS algorithms for dynamics are naturally resolved within tVMC. We demonstrate the power and generality of the method through four representative real-time problems in two dimensions: (I) chiral edge propagation in a free-fermion Chern insulator; (II) fractionalized charge transport in a fractional Chern insulator; (III) vison confinement dynamics in the Higgs phase of a $\mathbb{Z}_{2}$ lattice gauge theory coupled to a Higgs field; and (IV) superfluidity and critical velocity in interacting bosons. All simulations are performed on $12\times 12$ or $13\times 13$ lattices with evolution times $T=10$ to $12$ using modest computational resources — 1 to 5 days on a single GPU card. Where exact benchmarks exist (case I), PEPS–tVMC matches free-fermion dynamics with high accuracy up to $T=12$. These results establish PEPS–tVMC as a practical and versatile tool for real-time quantum dynamics in two dimensions. The method significantly extends the reach of classical tensor-network simulations for studying elementary excitations in quantum many-body systems and provides a valuable computational counterpart to emerging quantum simulators.
@@ I Introduction
Projected entangled pair states (PEPS) [19] provide a complementary variational family whose locality and multi-linearity make them particularly appealing for tVMC. Within VMC, only a single layer of PEPS tensors needs to be contracted, making the method substantially more efficient than conventional PEPS update schemes [20, 21, 22]. Furthermore, VMC allows parallel accumulation of samples, making the approach well suited for modern GPU architectures. In fact, PEPS-VMC has recently achieved considerable success for ground states, giving state-of-the-art results for the ground states of frustrated spins [23, 24, 25], fermion Hubbard model [26], and lattice gauge theory [27], owing to a foundational breakthrough in the sampling method of PEPS [28].
@@
While dynamical simulations with matrix product states (MPS) are enormously successful in 1D [33, 34, 35, 36], the situation in 2D is far more challenging. In contrast to the extensive 1D literature, only a small number of works—represented by Refs. [37, 38, 39, 40, 41, 42, 43, 44, 45, 46], have reported real-time simulations using PEPS or iPEPS, illustrating the general difficulty of the problem. Traditional approaches to dynamical PEPS have relied on full-update-like methods based on double-layer contraction [20, 21, 22], which face several challenges: (a) high cost of double layer contraction, (b) difficulty in fixing the PEPS gauge freedom, and (c) loss of positivity in the norm matrix approximately obtained from the contraction of the double-layer environment. We explain, in Sec. III.4, that the last two issues are naturally resolved while the first one is significantly alleviated within PEPS–tVMC, at the expense of manageable statistical noise.
@@ III Stochastic reconfiguration with PEPS
In this section we derive the SR equations for PEPS, analyze the structure of their null space, and describe how gauge redundancy can be eliminated efficiently in closed form, in both the SR and its reduced variant minSR [49] form.
Figure 1: PEPS structure and gauge freedom. (a) A $3\times 3$ PEPS network with local tensors AA. (b) A single tensor AA with physical index $p$ and virtual indices $(l,r,d,u)$. (c) Local gauge transformation on a virtual bond: inserting an invertible matrix $X$ and its inverse $X^{-1}$ between two adjacent tensors leaves the global PEPS wavefunction invariant. (d) Wavefunction amplitude $\braket{\bm{\mathbf{s}}|\Psi}$, computed via the contraction of the single layer network of sliced PEPS tensors: $a[\bm{\mathbf{x}}]_{lrdu}=A[\bm{\mathbf{x}}]^{p=\bm{\mathbf{s}}(\bm{\mathbf{x}})}_{lrdu}$.
@@
III.1 Gauge-fixing within SR
We now outline the procedure for eliminating null vectors within the SR formulation. First, we construct a projection matrix $P$ onto the subspace orthogonal to the constraint vectors $\bm{\mathbf{w}}$, thereby reducing the number of independent gauge vectors to
@@ III.1 Gauge-fixing within SR
III.2 Gauge-fixing within minSR
@@
We observe that the PEPS gauge redundancy does not appear at the minSR level, because the parameter index is contracted away, thus gauge-fixing is automatically achieved. Assuming decorrelation between samples, the matrix $\mathcal{O}^{\dagger}$ has a single null vector $1,1,\dots,1]^{T}$, which we remove with the same QR projection described above.
More generally, the null vector of $\mathcal{O}^{\dagger}$ is the weight vector $W$ used in performing the statistical average: $\braket{A}\approx\sum_{s=1}^{N_{s}}W_{s}A_{s}/N_{s}$, which in our case is simply the identity weight. A non-trivial weight vector would arise if reweighing is used in the Monte Carlo estimate. The removal of this null vector may be also benificial to NQS-tVMC in the minSR form.
Finally, we comment on the choice between SR and minSR. As the system size gets larger, $N_{p}$ eventually overtakes $N_{s}$ and one should use minSR for efficiency. However, it is always advised to study a problem first at small size and ideally benchmark against exact methods, in which case typically $N_{s}>N_{p}$ and it is only correct to use SR.
III.3 Small-$o$ trick in minSR
@@
from which the full minSR matrix $\mathcal{O}\mathcal{O}^{\dagger}$ can be reconstructed with the memory cost of $o$ instead of $O$ (see the Supplementary Material [50] for the explicit GPU-efficient construction).
@@
III.4 Comparison with double-layer methods
We briefly discuss the difficulties of double-layer methods mentioned in the introduction: computational cost, gauge fixing, and positivity of the norm matrix. Although manifested differently, they reflect the same underlying difficulty.
Computational cost.
Consider a PEPS with bond dimension $D$. Let $D^{\prime}$ denote the bond dimension of the boundary MPS used to contract a single-layer PEPS in VMC, cf. Fig. 1 (d), and likewise the boundary matrix product operator used in full-update double-layer contraction. The dominant computational costs scale as $D^{4}D^{\prime 2}$ for VMC [28] and $D^{4}D^{\prime 3}$ for full update [22]. Typical choices are $D^{\prime}\sim D$ to $D^{2}$ for VMC, and $D^{\prime}\sim D^{2}$ to $D^{3}$ for full update, leading to effective complexities of $D^{6}$$D^{8}$ and $D^{10}$$D^{13}$, respectively. This gap is substantial. VMC additionally incurs sampling overhead, but even choosing $D^{\prime}$ far larger than $D^{3}$ does not reliably stabilize the full-update procedure due to the next two issues.
@@ III.4 Comparison with double-layer methods
Unlike MPS in 1D [1], PEPS does not possess a canonical form that fixes the gauge freedoms. Ref. [22] introduces local gauge fixing for full-update gate applications, but no globally consistent gauge-fixing procedure is known. In fact, as an alternative, a variational subfamily — isometric PEPS (isoPEPS) [51, 41, 46] — is introduced to explicitly preserve the isometric gauge at the expense of representative expressiveness. However, the necessary shifting of isometric centers introduces significant errors and leads to suboptimal results. In contrast, PEPS–tVMC achieves global gauge fixing efficiently through a single QR projection or simply using the minSR formulation, without any loss of representational power.
@@
Because PEPS $\psi$ is an unnormalized ansatz, to perform a variational update, one needs to also form the norm matrix $N$ accounting for the norm $\braket{\psi|\psi}$. In practice, approximate contraction leads to violations of positivity of $N$, which destabilizes the generalized eigenvalue problems used in variational updates, especially in time evolution. Tiny errors in the boundary contraction can lead to large deviations in the time trajectory. tVMC avoids this issue entirely: normalization is enforced simply by subtracting $O_{\rm avg}$ and $h_{\rm avg}$ in Eq. 5, requiring no explicit norm matrix.
In the end, we mention that any conservation law in the computational basis can be implemented, in VMC, by sampling in the subsector of configurations with a definite total charge. This is equivalent to acting to the state a global projector of the desired charge sector, which significantly increases the representability of the underlying bare PEPS.
@@ IV Results
We now apply PEPS–tVMC to study four distinct two-dimensional models—bosonic and fermionic, free and interacting, and including systems with gauge symmetry or nontrivial topology. PEPS configurations are sampled using the sequential sampling algorithm, which was first introduced in Ref. [28], and later proved to satisfy detailed balance in Ref. [52]. The local energies and log-derivatives are computed via the contraction of the single-layer network shown in Fig. 1(d) (see Refs. [28, 52, 27] for methodological details). All simulation parameters are given in the Supplementary Material [50].
@@ IV.1 Free Fermion Chern insulator
As shown in Fig. 2, the excitation propagates unidirectionally along the edge. PEPS–tVMC faithfully reproduces the edge motion, its speed, and its spatial profile, in excellent agreement with the exact free-fermion numerics. For context, isoPEPS simulations of a similar chiral edge motion (e.g., Fig.6(b) of Ref.[46]) already provided state-of-the-art results at the time; however, their accuracy was ultimately limited by the errors due to the shifting of isometric centers. The present PEPS–tVMC approach achieves noticeably higher accuracy. We further note that the numerical difficulty of tensor network simulations is largely dictated by the smaller of $L_{x}$ and $L_{y}$. Thus, although the isoPEPS study used a $9\times 18$ lattice, the $12\times 12$ system considered here is in fact the more challenging geometry, making this benchmark particularly stringent.
Finally, we emphasize that the PEPS used in tVMC is entirely agnostic to the Gaussianity in the free fermion state. Thus, despite the model being noninteracting, its non-trivial dynamics makes a demanding test for PEPS, especially at this lattice size and time scale.
@@ IV.3 $\mathbb{Z}_{2}$ lattice gauge theory with Higgs field
Next, using the gauge-invariant PEPS–VMC framework recently introduced in Ref. [27], we examine the confinement dynamics in the $\mathbb{Z}_{2}$ gauge theory coupled to a Higgs field,
@@
subject to the Gauss-law constraint $\sigma_{i}^{z}\prod_{l\in i}Z_{l}=1$. $X_{l},Z_{l}$ are the Pauli $x,z$ matrices on the link $l$, representing the gauge field. $\sigma^{x}_{i},\sigma^{z}_{i}$ are the Pauli matrices on the site $i$, representing the Higgs field. $l_{\pm}$ are the sites at the ends of link $l$. Here $B_{p}=\prod_{l\in p}X_{l}$ measures the magnetic flux through plaquette $p$, and a violation of $B_{p}=1$ defines a vison excitation. The system has three ground state phases: confined (small $J$ and large $g$), Higgs (large $J$ and small $g$), and deconfined (small $J$ and small $g$) phases [64]. When $g=0$, the phase transition point between the deconfined phase and the Higgs phase is at $J_{c}\approx 0.328$ [65]. In the Higgs phase, the visons are expected to have a linearly attractive interaction, i.e. vison confinement. In the deconfined phase, the visons are expected to propagate freely, i.e. vison deconfinement.
@@
Recently, considerable progress has been made in implementing real-time gauge-theory dynamics on quantum devices [66, 67, 68]. Our results highlight that classical tensor-network approaches can serve as a highly effective and complementary platform for these studies. It would not be impossible for PEPS–tVMC to probe the intricate dynamics of glueballs in the (2+1)D Yang-Mills theory, after the gauge-invariant PEPS–VMC [27] is extended to the non-Abelian case.
@@ IV.4 Superfluid bosons
Figure 5: Real-time density response of a 2D bosonic superfluid to a moving obstacle (the red dot). Snapshots show the deviation of the local density from 0.5 on a $13\times 13$ lattice for different times $T$ as the obstacle is dragged along the $x$ direction. (a) $|\bm{\mathbf{v}}|=0.5$. (b) $|\bm{\mathbf{v}}|=5.0$.
@@ IV.5 Error in solving the SR equation
An important factor underlying the accuracy and efficiency of PEPS–tVMC is the numerical stability with which the SR equation—or its reduced minSR variant—can be solved. All simulations in this work solve the SR equation using a Cholesky decomposition [72] of the SR matrix with a diagonal regulator of magnitude $10^{-8}$. Compared to the rank-revealing singular value decomposition (SVD) solver, the Cholesky approach is substantially faster and more parallelizable on GPUs, but it lacks the robustness of SVD when applied to ill-conditioned matrices. Recent NQS–tVMC studies often resort to the SVD solver for this reason [13, 16, 18], trading computational efficiency for numerical robustness. The fact that Cholesky remains stable in all simulations indicates that once gauge redundancies are analytically removed, the PEPS variational manifold becomes exceptionally stable for real-time local quench dynamics.
@@ V Discussion and Outlook
The examples explored in Sec. IV demonstrate the breadth of physical phenomena accessible with this approach. These results establish PEPS–tVMC as a general-purpose tool for simulating elementary excitations, transport phenomena, and real-time response in a wide range of quantum systems with realistic computational resources. The improvements developed here apply equally to ground-state PEPS–VMC calculations [23, 24, 25, 26]. Extensions to excited states, finite temperature [73], and open quantum systems are also promising avenues. In the case of finite temperature, if purification-based ansatz is used, the small-$o$ trick can save the memory cost by a factor of 16 for spinful fermions!
@@ Acknowledgements.
Part of the algorithm is tested using the TenPy code base [74, 75]. The code for production is implemented with JAX [76]. Y.W. is supported by a startup grant from the IOP-CAS. I thank my colleagues Yu-Xiang Zhang, Shi-Xin Zhang, and Tao Xiang for sharing computational resources with me. I benefited greatly from discussion and collaboration on various related projects with Bartholomew Andrews, Zhehao Dai, Jannes Nys, Sheng-Hsuan Lin, Sajant Anand, Taige Wang, Michael Zaletel, Frank Pollmann, Wen-Yuan Liu, Jiahang Hu, Jiajun Yu, and Yuntian Gu.
@@
    2
    3def get_OOdag(o: jnp.ndarray, p: jnp.ndarray, K: int):
    4 r"""